OpenStudy (anonymous):

how could i integrate ( 5x/e^x ) ?

OpenStudy (anonymous):

\[\int\limits \frac{ 5x }{ e^{x} } dx = \int\limits 5x e^{-x} dx\] use partial integral > int u dv = uv + int v du

OpenStudy (anonymous):

where u = 5x , and dv = exp(-x) dx

OpenStudy (anonymous):

ok .. thanks alot

OpenStudy (anonymous):

ur welcome